OpenStudy (anonymous):

> >plsss help to my project a tennis ball whose initial speed is 2.0m/s rolls along a floor for 5.0m before coming to a top .find the coefficient of rolling friction. pls help me....^^

OpenStudy (roadjester):

Did the question specifically use the term "rolling friction"? http://hyperphysics.phy-astr.gsu.edu/hbase/frict2.html

OpenStudy (anonymous):

Initial kinetic energy (1/2) m v^2 will be dissipated by friction over a distance of 5 m This means that force x distance will be = original (1/2)m v^2 assuming the coefficient is defined as friction = c m g = c (weight) we have friction x distance = (c m g)(5) = (1/2)(m) v^2. c = (v^2) / 10 g where g is the acceleration of gravity, 9.8 m/s^2.